What concentration of Br results when 691 mL of 0.797 M KBr is mixed with 947 mL of 0.431 M FeBr,? [Br]= 0.750
Answer #1

This is mixing of the two salts. The complete calculation of the concentration is given as below:-

It is given that, volume of KBr = 691 mL = 0.691L. since, ImL = 10-3L. Concentration af KBr = 0.797 m =0.797 mol/L Number of

or Number of moles of RBn0 = 0.41 moles afen - © Total number of moles af en = (0,551 +0.41) molesolen number of moles af Bi